OpenStudy (love_to_love_you):

PLEASE HELP. FAN, MEDAL, AND TESTIMONY Suppose you are saving money for a new cell phone. At week 0 you have $20 and at week 10 you have $100. If y represents the amount of money you have saved and x represents the number of weeks, find and interpret the slope of the linear equation. a) 1/8 ; $8 is put away each week. b) 10; $10 is put away each week. c) 8; $8 is put away each week. d) 1/10; $10 is put away each week.

OpenStudy (love_to_love_you):

@kmeis002

OpenStudy (love_to_love_you):

@aum

OpenStudy (anonymous):

C, since you start with 20 and (8*10) equals 80, which added equals 100, and since slope is represented by\[y _{2}-y _{1} \div x _{2}-x _{1}\] 100-20 is 80 (y) and 10-0 is 10 (x), 80/10 simplifies to 8. Hope I helped!

OpenStudy (love_to_love_you):

Thank you so much!!! I understand it! Could you help me with a few more? @nickhunter172

OpenStudy (anonymous):

Yeah, sure. What you got for me?

OpenStudy (love_to_love_you):

Let me just type them real quick

OpenStudy (love_to_love_you):

You research the average cost of a tank of gasoline for a certain type of car for several recent years to look for trends. The table shows your data. What is the equation for a line of best fit? How much would you expect to pay for a tank of gas in the year 2019?

OpenStudy (anonymous):

May I have the table please?

OpenStudy (love_to_love_you):

a) y= 1.006x + 26.40; $47.52 b) y= 1.006x + 26.40; $58.92 c) y= 26.40x + 1.006; $47.52 d) y= 26.40x + 1.006; $58.92

OpenStudy (love_to_love_you):

@nickhunter172

OpenStudy (anonymous):

You can eliminate C and D straightaway because the slopes are ridiculous.

OpenStudy (love_to_love_you):

Yeah I was thinking it was a little weird to have that as a slope.

OpenStudy (love_to_love_you):

Is it A?

OpenStudy (anonymous):

The answer is A. Plug in 21 for x (0 is equivalent to 1998, since that is the first point given with the y-intercept 26.4, and 1998+21=2019.) 1.006*21= 21.126 [round down to 21.12 since you can't have 0.6 of a cent.] Add the y-intercept of 26.4, and get 47.52.

OpenStudy (love_to_love_you):

Yay I'm glad I at least got that right

OpenStudy (love_to_love_you):

If a function f(x) is shifted to the left one unit, what function represents the transformation?

OpenStudy (love_to_love_you):

a) f(x+1) b) f(x-1) c) f(x) + 1 d) f(x)-1

OpenStudy (love_to_love_you):

@nickhunter172

OpenStudy (anonymous):

I'm thinking B, but please graph it just to make sure.

OpenStudy (anonymous):

It is B, I was right. Constants outside of the bracket dictate vertical movement.

OpenStudy (love_to_love_you):

Okay, thats what I thought cause the graph was turning out to be that way. How many more questions can you handle? Haha I have a study guide and I am so unprepared for it due to being sick, and having a horrible teacher.

OpenStudy (anonymous):

Don't worry, my honors geometry teacher treats us like kids. I'll take any questions you got for me.

OpenStudy (love_to_love_you):

I'm taking Honors Algebra 2 and she is horrible. Thank you so much!

OpenStudy (love_to_love_you):

Let g(x) be the reflection of f(x)=x^2+5 in the x-axis. What is a function rule for g(x)? a) g(x) = -x^2 - 5 b) g(x) = x^2 + 5 c) g(x) = x^2 - 5 d) g(x) = -x^2 + 5

OpenStudy (anonymous):

-5 changes the y-intercept of the function, so C would be the most accurate reflection.

OpenStudy (love_to_love_you):

Okay

OpenStudy (love_to_love_you):

What is the equation of the absolute value function?

OpenStudy (love_to_love_you):

a) y= 6|x+2|-2 b) y= 6|-x-2|+2 c) y= -6|x+2|-2 d) y= -6|x-2|-2

OpenStudy (love_to_love_you):

I think it's b

OpenStudy (anonymous):

Yep, that would be it.

OpenStudy (love_to_love_you):

Okay so I have 3 more questions and they require that I show all work. Every step. Can you help me on those or should I open a new question?

OpenStudy (love_to_love_you):

@nickhunter172

OpenStudy (love_to_love_you):

I wish you could help me with my portfolio and test, you're so amazing. Hahaha

OpenStudy (anonymous):

I can help, and thank you. :3